# Geometric control of boundary-catalytic branching processes

### v1

@@ Abstract
Boundary-catalytic branching processes describe a broad class of natural phenomena where the population of diffusing particles grows due to their spontaneous binary branching (e.g., division, fission or splitting) on a catalytic boundary located in a complex environment. We investigate the possibility of geometric control of the population growth by compensating the proliferation of particles due to branching events by their absorptions in the bulk or on boundary absorbing regions. We identify an appropriate Steklov spectral problem to obtain the phase diagram of this out-of-equilibrium stochastic process. The principal eigenvalue determines the critical line that separates an exponential growth of the population from its extinction. In other words, we establish a powerful tool for calculating the optimal absorption rate that equilibrates the opposite effects of branching and absorption events and thus results in steady-state behavior of this diffusion-reaction system. Moreover, we show the existence of a critical catalytic rate above which no compensation is possible, so that the population cannot be controlled and keeps growing exponentially. The proposed framework opens promising perspectives for better understanding, modeling and control of various boundary-catalytic branching processes, with applications in physics and life sciences.
@@
Branching processes are critically important in disciplines as diverse as atomic energy (neutron production in a nuclear reactor), optics (stimulated emission in a laser cavity), biology (bacterial colony growth), ecology (population dynamics), social sciences (genealogy), to name but a few [Harris, Williams, Kimmel]. In the simplest homogeneous setting, each particle can spontaneously split into two identical copies that undergo the same branching process independently of each other. These stochastic branching events result in an exponential growth of the population size in time. In most applications in Nature and industry, however, branching processes are spatially heterogeneous and occur on catalytic sites [Dawson99, Klenke00, Kesten03, Bulinskaya18]. Such catalytic branching processes strongly rely on the motion of particles and thus couple branching mechanisms to particle dynamics. For instance, when a bacterial colony explores the available space foraging for food, the population growth is controlled by available resources, their amount and spatial configuration.
@@
In this Letter, we consider the so-called boundary-catalytic branching (BCB) processes when particles diffuse in a restricted environment and branching events occur at boundaries: physical, spatial, or functional. This boundary-driven behavior appears in many real-world settings. In physics, such models describe particle systems where reactions occur only at interfaces, such as catalytic surfaces in chemical kinetics [Ben-Avraham, Lindenberg, Galanti16, Grebenkov23n]. In life sciences, they help explaining phenomena like cell proliferation at tissue boundaries, stem-cell niche dynamics, the spread of populations or epidemics when reproduction or contamination is concentrated at ecological edges [Murrey, Schuss, Bressloff13]. By combining stochastic branching with spatial constraints, boundary-catalytic models reveal how local interactions at borders can shape global system behavior, making them valuable tools for understanding complex, heterogeneous environments across scientific disciplines.
@@
Being inspired by applications in life sciences, we investigate a critically relevant possibility of geometric control of the population size: how can one incorporate absorbing regions into the environment to efficiently eliminate excessive particles in order to put an explosive population growth under control? For this purpose, we employ various tools such as a probabilistic construction of the BCB process based on the boundary local time, a spectral representation of the population size and its long-time asymptotic behavior, matched asymptotic expansions, and optimization schemes for choosing the appropriate reactivity to balance branching and absorption events. Despite intensive research on diffusion-mediated phenomena and related first-passage statistics [Redner, Krapivsky, Metzler, Grebenkov, Bray13, Benichou14, Levernier19], the important topic of BCB processes is yet unexplored in physics literature. In fact, even though the general concept of balancing birth and death events is very natural and has been studied in the past (see, e.g., [Canet04a, Canet04b, Odor04] for the analysis of phase diagrams for branching-annihilating random walks), the spatial location of branching and absorption events on a boundary brings a new dimension to this fundamental problem.
@@
We consider the following model of BCB processes. At time 0, a single particle is released from a point $\bm{x}$ and diffuses inside a confining domain $\Omega\subset{\mathbb{R}}^{d}$ with a constant diffusivity $D$. The boundary of this environment, $\partial\Omega=\Gamma_{\mathrm{c}}\cup\Gamma_{\mathrm{a}}\cup\Gamma_{\mathrm{r}}$, is in general partitioned into three disjoint subsets: a catalytic region $\Gamma_{\mathrm{c}}$, an absorbing region $\Gamma_{\mathrm{a}}$, and the remaining reflecting region $\Gamma_{\mathrm{r}}$. As illustrated on Fig. 1(a), both catalytic and absorbing regions can be composed of multiple “pieces”, either located on the boundary or in the bulk. The absorbing region is characterized by a reactivity $\kappa_{\mathrm{a}}=q_{\mathrm{a}}D>0$, which can range from 0 (no absorption) to $\infty$ (a perfect sink that instantly kills the particle upon its first arrival onto $\Gamma_{\mathrm{a}}$). When $q_{\mathrm{a}}$ is finite, each particle arrived onto the absorbing region $\Gamma_{\mathrm{a}}$ is either absorbed (with a small probability proportional to $q_{\mathrm{a}}$), or reflected back to resume its diffusion in $\Omega$ [Grebenkov03, Erban07, Singer08, Piazza22]. As a result, the particle is killed when the random number of its encounters with the absorbing region, represented by the boundary local time $\ell_{t}^{\mathrm{a}}$ on $\Gamma_{\mathrm{a}}$, exceeds an independent random threshold $\hat{\ell}^{\mathrm{a}}$ obeying the exponential distribution ${\mathbb{P}}\{\hat{\ell}^{\mathrm{a}}>\ell\}=e^{-q_{\mathrm{a}}\ell}$ [Grebenkov20]. The parameter $q_{\mathrm{a}}$ quantifies thus the intensity of absorption events. With some abuse of language, we will call $q_{\mathrm{a}}$ the absorption rate, even though $q_{\mathrm{a}}$ has units of the inverse of length, in accordance to the fact that the boundary local time $\ell_{t}^{\mathrm{a}}$ has units of length. A similar construction was employed in the probabilistic description of permeation processes such as snapping out Brownian motion [Lejay16, Bressloff22, Bressloff23].
@@
In the same vein, the catalytic region $\Gamma_{\mathrm{c}}$ is characterized by a reactivity $\kappa_{\mathrm{c}}=q_{\mathrm{c}}D\geq 0$, so that the particle hitting $\Gamma_{\mathrm{c}}$ can either be replaced by two new particles (with a small probability proportional to $q_{\mathrm{c}}$), or be reflected back to resume its diffusion (Fig. 1(b)). In other words, the branching event is triggered when the boundary local time $\ell_{t}^{\mathrm{c}}$ of the particle on $\Gamma_{\mathrm{c}}$ exceeds another independent random threshold $\hat{\ell}^{\mathrm{c}}$ obeying the exponential distribution ${\mathbb{P}}\{\hat{\ell}^{\mathrm{c}}>\ell\}=e^{-q_{\mathrm{c}}\ell}$, with the catalytic rate (or branching intensity) $q_{\mathrm{c}}$. The two newborn particles are released from the position of their branching event (with their boundary local times set to 0) and diffuse independently of each other, until the next branching or absorption event, and so on (see [DelGross76, Delmas05, Bocharov14] for further mathematical details). The competition between the opposite effects of catalytic and absorbing regions will determine the population dynamics and, in particular, the population size $N(t|\bm{x})$ (i.e., the mean number of particles at time $t$). Even though the above probabilistic constructions of absorption and branching events are identical, we will show that the impact of the rates $q_{\mathrm{a}}$ and $q_{\mathrm{c}}$ onto the population size is dramatically different.
@@
As the number of diffusing particles and thus the dimension of the configuration space randomly change in time, a probabilistic description of branching processes requires elaborate mathematical tools such as measure-valued stochastic processes (or superprocesses) and nonlinear PDEs [LeGall, Dynkin, Englander04, Morters05, Englander07]. In turn, the population size $N(t|\bm{x})$ still admits the linear PDE description via the diffusion equation with Robin boundary condition
@@
with the initial condition $N(0|\bm{x})=1$. Here $\partial_{n}$ is the normal derivative oriented outward the domain $\Omega$, $\Delta$ is the Laplace operator governing the diffusion dynamics, and $q(\bm{x})$ is the piecewise-constant function: $q(\bm{x})=q_{\mathrm{a}}$ on the absorbing region $\Gamma_{\mathrm{a}}$, $q(\bm{x})=0$ on the reflecting region $\Gamma_{\mathrm{r}}$, and $q(\bm{x})=-q_{\mathrm{c}}$ on the catalytic region $\Gamma_{\mathrm{c}}$. The negative parameter $q_{\mathrm{c}}$ in the Robin boundary condition incorporates the catalytic effect of $\Gamma_{\mathrm{c}}$, which plays the role of a source of particles that are released into the bulk after branching. In fact, the left-hand side of Eq. (1b) is the net diffusive flux to the bulk, which is set to be equal to the flux of created particles, the latter being proportional to $N(t|\bm{x})$. The presence of the negative parameter $q_{\mathrm{c}}$ makes the crucial difference from the conventional analysis of the survival probability that admits the same PDE and characterizes first-passage times, target search problems, and diffusion-controlled reactions in reactive media [Redner, Krapivsky, Murrey, Schuss, Metzler, Lindenberg, Galanti16, Grebenkov, Bray13, Bressloff13, Benichou14, Levernier19, Grebenkov23n]. In this study, we aim at answering how the geometric configuration of absorbing and catalytic regions controls the long-time behavior of the population size $N(t|\bm{x})$. We will treat separately the cases of bounded and unbounded environments.
@@
The eigenvalues are enumerated by the index $k=0,1,2,\cdots$ to form an increasing sequence, $\lambda_{0}\leq\lambda_{1}\leq\cdots\nearrow+\infty$, whereas the eigenfunctions $\{u_{k}\}$ form a complete orthonormal basis of $L^{2}(\Omega)$ [Levitin]. As a consequence, the solution of the diffusion equation (1) admits a standard spectral expansion,
@@
It is therefore clear that the long-time behavior of $N(t|\bm{x})$ is controlled by the principal (smallest) eigenvalue $\lambda_{0}$: $N(t|\bm{x})\propto e^{-Dt\lambda_{0}}$ as $t\to\infty$. In sharp contrast to the survival probability, which always decays exponentially due to eventual absorption on $\Gamma_{\mathrm{a}}$ with $\kappa_{\mathrm{a}}>0$, the behavior of the population size depends on the competition between the proliferative effect of branching events on $\Gamma_{\mathrm{c}}$ and the destroying effect of absorption events on $\Gamma_{\mathrm{a}}$. In other words, $N(t|\bm{x})$ may exponentially decay ($\lambda_{0}>0$), reach a steady-state limit ($\lambda_{0}=0$), or exponentially grow ($\lambda_{0}<0$). The geometric control of the population growth has therefore two aspects: (i) for a given geometric setting, to find the optimal absorption rate $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},\lambda_{0})$ that yields the desired value of $\lambda_{0}$; (ii) for given $q_{\mathrm{a}}$ and $q_{\mathrm{c}}$, to find the optimal geometric configuration of absorption regions to achieve the desired $\lambda_{0}$. In the following, we will mainly focus on the first aspect for the target value $\lambda_{0}=0$ ensuring the steady-state limit and thus separating the exponential growth from extinction on the phase diagram in the $(q_{\mathrm{c}},q_{\mathrm{a}})$ space.
In the low-rate regime when both $q_{\mathrm{c}}$ and $q_{\mathrm{a}}$ are small, Robin boundary conditions in Eqs. (2) are close to the Neumann one so that the principal eigenvalue $\lambda_{0}$ is close to zero, whereas the associated eigenfunction $u_{0}$ is close to a constant. A standard perturbation approach yields in the leading order (see Supplementary Material SM.I for details)
@@
In the following, we investigate geometric control beyond this approximate relation and reveal its limitations.
To gain some intuition onto this geometric control, let us first inspect the case of an interval of length $L$, $\Omega=(0,L)$, with $\Gamma_{\mathrm{c}}=\{0\}$, $\Gamma_{\mathrm{a}}=\{L\}$ and $\Gamma_{\mathrm{r}}=\emptyset$ (no reflecting region). The Laplacian eigenfunctions are obtained as linear combinations of sine and cosine functions in a standard way (see SM.II for details); in particular, the eigenvalues $\lambda_{k}$ are the solutions of the transcendental equation:
@@
When $q_{\mathrm{c}}=0$ (no branching, just reflections at $x=0$), one retrieves the classical problem of the survival of a diffusing particle on the interval with one absorbing and one reflecting endpoint [Carslaw]. In turn, if $q_{\mathrm{c}}=q_{\mathrm{a}}$, Eq. (6) has infinitely many positive solutions $\lambda_{k}=\pi^{2}k^{2}/L^{2}$, with $k=1,2,\cdots$, as well as one negative solution: $\lambda_{0}=-q_{\mathrm{a}}q_{\mathrm{c}}$. One sees that, even though branching and absorption events are governed by the same mechanism, the catalytic effect always “wins” in this setting, yielding $\lambda_{0}<0$. In order to balance the catalytic effect, one therefore needs absorption events at higher rate. Setting $\lambda_{k}=0$ in Eq. (6) yields $q_{\mathrm{a}}q_{\mathrm{c}}L=q_{\mathrm{a}}-q_{\mathrm{c}}$ so that the steady-state regime can only emerge at the optimal absorption rate $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)=q_{\mathrm{c}}/(1-q_{\mathrm{c}}L)$ and only if $q_{\mathrm{c}}\leq q_{\mathrm{c}}^{\rm crit}=1/L$. In fact, if $q_{\mathrm{c}}$ exceeds the critical value $q_{\mathrm{c}}^{\rm crit}$, even the perfect sink with $q_{\mathrm{a}}=\infty$ cannot compensate the catalytic effect, implying the exponential growth of the population size.
The above explicit analysis can be extended to higher dimensions $d\geq 2$ by considering diffusion in a spherical shell between a catalytic surface of radius $R$ with $q_{\mathrm{c}}>0$ and a concentric absorbing surface of radius $L$ with $q_{\mathrm{a}}\geq 0$. Solving Eqs. (2) with $\lambda=0$ via the separation of variables, we get (see SM.III for details):
@@
Most importantly, the above intuitive picture remains valid for general bounded domains. The eigenvalue problem (2) that we used to determine the principal eigenvalue $\lambda_{0}$, can also be regarded as a mathematical tool for the geometric control of BCB processes. In fact, fixing the catalytic rate $q_{\mathrm{c}}$ and the desired value of $\lambda$, one can treat $q_{\mathrm{a}}$ as a spectral parameter, denoted as $\sigma$, and search for eigenpairs $\{\sigma,u\}$ satisfying Eqs. (2). This is the so-called generalized Steklov spectral problem, which is known to have a discrete spectrum [Levitin]. Due to the presence of the catalytic region $\Gamma_{\mathrm{c}}$ with the negative Robin parameter $q_{\mathrm{c}}$, the principal (smallest) eigenvalue $\sigma_{0}$ is strictly negative. This eigenvalue determines the optimal absorption rate, $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},\lambda_{0})=-\sigma_{0}$, that one has to use to achieve the desired value $\lambda_{0}$ in a given geometric setting. Moreover, the associated eigenfunction $u_{0}(\bm{x})$ determines the steady-state spatial profile of the population size $N(\infty|\bm{x})$. Alternatively, one could fix $q_{\mathrm{a}}$ and treat $q_{\mathrm{c}}$ as a spectral parameter $\mu$ to determine the dependence of $q_{\mathrm{c}}$ on $\lambda_{0}$ and $q_{\mathrm{a}}$. For instance, if we are interested in the steady-state solution ($\lambda_{0}=0$) with the maximal absorption rate $q_{\mathrm{a}}=\infty$ (a perfect sink), Eq. (2) is reduced to

### v2

# The geometric control of boundary-catalytic branching processes

@@ Abstract
Boundary-catalytic branching processes describe a broad class of natural phenomena where the population of diffusing particles grows due to their spontaneous binary branching (e.g., division, fission or splitting) on a catalytic boundary located in a complex environment. We investigate the possibility of the geometric control of the population growth by compensating the proliferation of particles due to catalytic branching events by their absorptions in the bulk or on absorbing regions of the boundary. We identify an appropriate Steklov spectral problem to obtain the phase diagram of this out-of-equilibrium stochastic process. The principal eigenvalue determines the critical line that separates an exponential growth of the population from its extinction in a bounded domain. In other words, we establish a powerful tool for calculating the optimal absorption rate that equilibrates the opposite effects of branching and absorption events and thus results in steady-state behavior of this diffusion-reaction system. Moreover, we show the existence of a critical catalytic rate above which no compensation is possible, so that the population cannot be controlled and keeps growing exponentially. The proposed framework opens promising perspectives for better understanding, modeling and control of various boundary-catalytic branching processes, with applications in physics, chemistry, and life sciences.
@@ I Introduction
Branching processes are critically important in disciplines as diverse as atomic energy (neutron production in a nuclear reactor), optics (stimulated emission in a laser cavity), chemistry (heterogeneous catalysis), biology (bacterial colony growth), ecology (population dynamics), social sciences (genealogy), to name but a few [1, 2, 3]. In the simplest homogeneous setting, each particle can spontaneously split into two identical copies that undergo the same branching process independently of each other. These stochastic branching events result in an exponential growth of the population size in time. In most applications in Nature and industry, however, branching processes are spatially heterogeneous and occur on catalytic sites [4, 5, 6, 7]. Such catalytic branching processes strongly rely on the motion of particles and thus couple branching mechanisms to particle dynamics. For instance, when a bacterial colony explores the available space foraging for food, the population growth is controlled by available resources, their amount and spatial configuration.
@@
In this paper, we consider the so-called boundary-catalytic branching (BCB) processes when particles diffuse in a restricted environment and branching events occur at boundaries: physical, spatial, or functional. This boundary-driven behavior appears in many real-world settings. In physics, such models describe particle systems where reactions occur only at interfaces, such as catalytic surfaces in chemical kinetics [8, 9, 10, 11]. In life sciences, they help explaining phenomena like cell proliferation at tissue boundaries, stem-cell niche dynamics, the spread of populations or epidemics when reproduction or contamination is concentrated at ecological edges [12, 13, 14]. By combining stochastic branching with spatial constraints, boundary-catalytic models reveal how local interactions at borders can shape global system behavior, making them valuable tools for understanding complex, heterogeneous environments across scientific disciplines.
@@
Inspired by applications in life sciences, we investigate a critically relevant possibility of the geometric control of the population size: how can one incorporate absorbing regions into the environment to efficiently eliminate excessive particles in order to put an explosive population growth under control? For this purpose, we employ various tools such as a probabilistic construction of the BCB process based on the boundary local time, a spectral representation of the mean population size and its long-time asymptotic behavior, matched asymptotic expansions, and optimization schemes for choosing the appropriate reactivity to balance branching and absorption events. Despite intensive research on diffusion-mediated phenomena and related first-passage statistics [15, 16, 17, 18, 19, 20, 21], the important topic of BCB processes is yet unexplored in physics literature. In fact, even though the general concept of balancing birth and death events is very natural and has been studied in the past (see, e.g., [22, 23, 24] for the analysis of phase diagrams for branching-annihilating random walks), the spatial location of branching and absorption events on a boundary brings a new dimension to this fundamental problem.
II Model
@@
We consider the following model of BCB processes. At time 0, a single particle is released from a point $\bm{x}$ and diffuses inside a confining domain $\Omega\subset{\mathbb{R}}^{d}$ with a constant diffusivity $D$. The boundary of this environment, $\partial\Omega=\Gamma_{\mathrm{c}}\cup\Gamma_{\mathrm{a}}\cup\Gamma_{\mathrm{r}}$, is in general partitioned into three disjoint subsets: a catalytic region $\Gamma_{\mathrm{c}}$, an absorbing region $\Gamma_{\mathrm{a}}$, and the remaining reflecting region $\Gamma_{\mathrm{r}}$. As illustrated on Fig. 1(a), both catalytic and absorbing regions can be composed of multiple “pieces”, either located on the boundary or in the bulk. The absorbing region is characterized by a reactivity $\kappa_{\mathrm{a}}=q_{\mathrm{a}}D\geq 0$, which can range from 0 (no absorption) to $\infty$ (a perfect sink that instantly kills the particle upon its first arrival onto $\Gamma_{\mathrm{a}}$). When $q_{\mathrm{a}}$ is finite, each particle arrived onto the absorbing region $\Gamma_{\mathrm{a}}$ is either absorbed (with a small probability proportional to $q_{\mathrm{a}}$), or reflected back to resume its diffusion in $\Omega$ [25, 26, 27, 28]. As a result, the particle is killed when the random number of its encounters with the absorbing region, represented by the boundary local time $\ell_{t}^{\mathrm{a}}$ on $\Gamma_{\mathrm{a}}$, exceeds an independent random threshold $\hat{\ell}^{\mathrm{a}}$ obeying the exponential distribution ${\mathbb{P}}\{\hat{\ell}^{\mathrm{a}}>\ell\}=e^{-q_{\mathrm{a}}\ell}$ [29]. The parameter $q_{\mathrm{a}}$ quantifies thus the intensity of absorption events. With some abuse of language, we will call $q_{\mathrm{a}}$ the absorption rate, even though $q_{\mathrm{a}}$ has units of the inverse of length, in accordance to the fact that the boundary local time $\ell_{t}^{\mathrm{a}}$ has units of length. A similar construction was employed in the probabilistic description of permeation processes such as snapping out Brownian motion [30, 31, 32].
@@
In the same vein, the catalytic region $\Gamma_{\mathrm{c}}$ is characterized by a reactivity $\kappa_{\mathrm{c}}=q_{\mathrm{c}}D\geq 0$, so that the particle hitting $\Gamma_{\mathrm{c}}$ can either be replaced by two new particles (with a small probability proportional to $q_{\mathrm{c}}$), or be reflected back to resume its diffusion (Fig. 1(b)). In other words, the branching event is triggered when the boundary local time $\ell_{t}^{\mathrm{c}}$ of the particle on $\Gamma_{\mathrm{c}}$ exceeds another independent random threshold $\hat{\ell}^{\mathrm{c}}$ obeying the exponential distribution ${\mathbb{P}}\{\hat{\ell}^{\mathrm{c}}>\ell\}=e^{-q_{\mathrm{c}}\ell}$, with the catalytic rate (or branching intensity) $q_{\mathrm{c}}$. The two newborn particles are released from the position of their branching event (with their boundary local times set to 0) and diffuse independently of each other, until the next branching or absorption event, and so on (see [33, 34, 35] for further mathematical details). The competition between the opposite effects of catalytic and absorbing regions will determine the population dynamics and, in particular, the mean population size $N(t|\bm{x})$ (i.e., the average number of particles at time $t$). Even though the above probabilistic constructions of absorption and branching events are identical, we will show that the impact of the rates $q_{\mathrm{a}}$ and $q_{\mathrm{c}}$ onto the population size is dramatically different.
@@
III Results
III.1 Negative reactivity on the catalytic region
As the number of diffusing particles and thus the dimension of the configuration space randomly change in time, a probabilistic description of branching processes requires elaborate mathematical tools such as measure-valued stochastic processes (or superprocesses) and nonlinear PDEs [36, 37, 38, 39, 40]. In turn, the mean population size $N(t|\bm{x})$ still admits the linear PDE description via the diffusion equation with Robin boundary condition,
@@ III.1 Negative reactivity on the catalytic region
with the initial condition $N(0|\bm{x})=1$. Here $\partial_{n}$ is the normal derivative oriented outward the domain $\Omega$, $\Delta$ is the Laplace operator governing the diffusion dynamics, and $q(\bm{x})$ is the piecewise-constant function: $q(\bm{x})=q_{\mathrm{a}}$ on the absorbing region $\Gamma_{\mathrm{a}}$, $q(\bm{x})=0$ on the reflecting region $\Gamma_{\mathrm{r}}$, and $q(\bm{x})=-q_{\mathrm{c}}$ on the catalytic region $\Gamma_{\mathrm{c}}$. The negative parameter $q_{\mathrm{c}}$ in the Robin boundary condition incorporates the catalytic effect of $\Gamma_{\mathrm{c}}$, which plays the role of a source of particles that are released into the bulk after branching. In fact, the left-hand side of Eq. (1b) is the net diffusive flux to the bulk, which is set to be equal to the flux of created particles, the latter being proportional to $N(t|\bm{x})$. The PDE description (1) in the purely catalytic case (i.e., $\partial\Omega=\Gamma_{\mathrm{c}}$) was rigorously established in [33] for polygonal bounded domains but its extension to our mixed setting with a piecewise constant rate $q(\bm{x})$, other domains and higher dimensions seems to be straightforward. Moreover, as branching and absorption events occur exclusively on the boundary, we expect that ordinary diffusion can be replaced by a general Markov process governed by a second-order elliptic operator, in which case the diffusion equation (1a) has to be substituted by the backward Fokker-Planck (or Kolmogorov) equation. A mathematical justification of such an extension and the analysis of its consequences are beyond the scope of this paper.
The presence of the negative parameter $q_{\mathrm{c}}$ makes a crucial difference from the conventional analysis of the survival probability that admits the same PDE description and characterizes first-passage times, target search problems, and diffusion-controlled reactions in reactive media [15, 16, 12, 13, 17, 9, 10, 18, 19, 14, 20, 21, 11]. In this study, we aim at answering how the geometric configuration of absorbing and catalytic regions controls the long-time behavior of the mean population size $N(t|\bm{x})$. We will treat separately the cases of bounded and unbounded environments.
III.2 The geometric control of the population growth
@@
The eigenvalues are enumerated by the index $k=0,1,2,\cdots$ to form an increasing sequence, $\lambda_{0}\leq\lambda_{1}\leq\cdots\nearrow+\infty$, whereas the eigenfunctions $\{u_{k}\}$ form a complete orthonormal basis of $L^{2}(\Omega)$ [41]. As a consequence, the solution of the diffusion equation (1) admits a standard spectral expansion,
@@
It is therefore clear that the long-time behavior of $N(t|\bm{x})$ is controlled by the principal (smallest) eigenvalue $\lambda_{0}$: $N(t|\bm{x})\propto e^{-Dt\lambda_{0}}$ as $t\to\infty$. In sharp contrast to the survival probability, which always decays exponentially due to eventual absorption on $\Gamma_{\mathrm{a}}$ with $\kappa_{\mathrm{a}}>0$, the behavior of the mean population size depends on the competition between the proliferative effect of branching events on $\Gamma_{\mathrm{c}}$ and the destroying effect of absorption events on $\Gamma_{\mathrm{a}}$. In other words, $N(t|\bm{x})$ may exponentially decay ($\lambda_{0}>0$), reach a steady-state limit ($\lambda_{0}=0$), or exponentially grow ($\lambda_{0}<0$). The geometric control of the population growth has therefore two aspects: (i) for a given geometric setting, to find the optimal absorption rate $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},\lambda_{0})$ that yields the desired value of $\lambda_{0}$ for a given catalytic rate $q_{\mathrm{c}}$; (ii) for given $q_{\mathrm{a}}$ and $q_{\mathrm{c}}$, to find the optimal geometric configuration of absorption regions to achieve the desired $\lambda_{0}$. In the following, we will mainly focus on the first aspect for the target value $\lambda_{0}=0$ ensuring the steady-state limit and thus separating the exponential growth from extinction on the phase diagram in the $(q_{\mathrm{c}},q_{\mathrm{a}})$ space.
In the low-rate regime when both $q_{\mathrm{c}}$ and $q_{\mathrm{a}}$ are small, Robin boundary conditions in Eqs. (2) are close to the Neumann one so that the principal eigenvalue $\lambda_{0}$ is close to zero, whereas the associated eigenfunction $u_{0}$ is close to a constant. A standard perturbation approach yields in the leading order (see Appendix A for details)
@@
In the following, we investigate the geometric control beyond this approximate relation and reveal its limitations.
III.3 Explicit examples
To gain some intuition onto the geometric control, let us first inspect the case of an interval of length $L$, $\Omega=(0,L)$, with $\Gamma_{\mathrm{c}}=\{0\}$, $\Gamma_{\mathrm{a}}=\{L\}$ and $\Gamma_{\mathrm{r}}=\emptyset$ (no reflecting region). The Laplacian eigenfunctions are obtained as linear combinations of sine and cosine functions in a standard way (see Appendix B for details); in particular, the eigenvalues $\lambda_{k}$ are the solutions of the transcendental equation:
@@ III.3 Explicit examples
When $q_{\mathrm{c}}=0$ (no branching, just reflections at $x=0$), one retrieves the classical problem of the survival of a diffusing particle on the interval with one absorbing and one reflecting endpoint [42]. In turn, if $q_{\mathrm{c}}=q_{\mathrm{a}}$, Eq. (6) has infinitely many positive solutions $\lambda_{k}=\pi^{2}k^{2}/L^{2}$, with $k=1,2,\cdots$, as well as one negative solution: $\lambda_{0}=-q_{\mathrm{a}}q_{\mathrm{c}}$. One sees that, even though branching and absorption events are governed by the same mechanism, the catalytic effect always “wins” in this setting, yielding $\lambda_{0}<0$. In order to balance the catalytic effect, one therefore needs absorption events at higher rate. Setting $\lambda_{k}=0$ in Eq. (6) yields $q_{\mathrm{a}}q_{\mathrm{c}}L=q_{\mathrm{a}}-q_{\mathrm{c}}$ so that the steady-state regime can only emerge at the optimal absorption rate $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)=q_{\mathrm{c}}/(1-q_{\mathrm{c}}L)$ and only if $q_{\mathrm{c}}\leq q_{\mathrm{c}}^{\rm crit}=1/L$. In fact, if $q_{\mathrm{c}}$ exceeds the critical value $q_{\mathrm{c}}^{\rm crit}$, even the perfect sink with $q_{\mathrm{a}}=\infty$ cannot compensate the catalytic effect, implying the exponential growth of the mean population size.
The above explicit analysis can be extended to higher dimensions $d\geq 2$ by considering diffusion in a spherical shell $\Omega=\{\bm{x}\in{\mathbb{R}}^{d}~:~R<|\bm{x}|<L\}$ between a catalytic sphere of radius $R$ with $q_{\mathrm{c}}>0$ and a concentric absorbing sphere of radius $L$ with $q_{\mathrm{a}}\geq 0$. Solving Eqs. (2) with $\lambda=0$ via the separation of variables, we get (see Appendix C for details):
@@
III.4 Phase diagram for bounded domains
Most importantly, the above intuitive picture remains valid for general bounded domains. The eigenvalue problem (2) that we used to determine the principal eigenvalue $\lambda_{0}$, can also be regarded as a mathematical tool for the geometric control of BCB processes. In fact, fixing the catalytic rate $q_{\mathrm{c}}$ and the desired value of $\lambda$, one can treat $q_{\mathrm{a}}$ as a spectral parameter, denoted as $\sigma$, and search for eigenpairs $\{\sigma,u\}$ satisfying Eqs. (2). This is the so-called generalized Steklov spectral problem, which is known to have a discrete spectrum [41]. Due to the presence of the catalytic region $\Gamma_{\mathrm{c}}$ with the negative Robin parameter $q_{\mathrm{c}}$, the principal (smallest) eigenvalue $\sigma_{0}$ is strictly negative. This eigenvalue determines the optimal absorption rate, $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},\lambda_{0})=-\sigma_{0}$, that one has to use to achieve the desired value $\lambda_{0}$ in a given geometric setting. In the particular case $\lambda_{0}=0$, the function $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ determines the critical line in the $(q_{\mathrm{c}},q_{\mathrm{a}})$ plane that separates the extinction regime ($q_{\mathrm{a}}>\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$) from the growth regime ($q_{\mathrm{a}}<\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$) on the phase diagram. Moreover, the associated eigenfunction $u_{0}(\bm{x})$ determines the steady-state spatial profile of the mean population size $N(\infty|\bm{x})$ (further spectral insights are discussed in Appendix D). Alternatively, one could fix $q_{\mathrm{a}}$ and treat $q_{\mathrm{c}}$ as a spectral parameter $\mu$ to determine the function $\hat{q}_{\mathrm{c}}(q_{\mathrm{a}},\lambda_{0})$ that fixes $q_{\mathrm{c}}$ to achieve the desired value of $\lambda_{0}$ for a given $q_{\mathrm{a}}$. For instance, if we are interested in the steady-state solution ($\lambda_{0}=0$) with the maximal absorption rate $q_{\mathrm{a}}=\infty$ (a perfect sink), Eq. (2) is reduced to